Prove that every ideal I of a ring R is the kernel of a ring homomorphism of R.


1
Expert's answer
2021-06-18T02:29:01-0400

Let us prove that every ideal "I" of a ring "R" is the kernel of a ring homomorphism of "R". Consider the quotient ring "R/I=\{a+I:a\in R\}." Let "\overline{a}" denotes "a+I." Then "\overline{a}+\overline{b}=\overline{a+b},\ \overline{a}\cdot\overline{b}=\overline{a\cdot b}." Consider the canonical ring homomorphism "\pi:R\to R/I,\ \pi(a)=\overline{a}." It is indeed a homomorphism because of "\pi(a+b)=\overline{a+b}=\overline{a}+\overline{b}=\pi(a)+\pi(b)," "\pi(a\cdot b)=\overline{a\cdot b}=\overline{a}\cdot \overline{b}=\pi(a)\cdot\pi(b)." It follows that "\ker\pi=\{a\in R: \pi(a)=\overline{0}\}=\{a\in R: \overline{a}=\overline{0}\}
=\{a\in R: a+I=0+I\}=\{a\in R: a\in I-I\}=\{a\in R: a\in I\}=I."

We conclude that "I" is a kernel of a ring homomorphism "\pi:R\to R/I."